OpenStudy (anonymous):

18 is 12% of what number?

OpenStudy (zzr0ck3r):

.12x=18

OpenStudy (anonymous):

If I'm not wrong, its 150... 150 * 0.12 = 18

OpenStudy (xxdragoonxx):

18/12=1.5

OpenStudy (anonymous):

No divide .12 by 18 and that is your answer!

OpenStudy (anonymous):

I Think!

OpenStudy (zzr0ck3r):

nope 18 is 12% of what number 18 = .12 x 18=0.12*x x=18/0.12 = 150